Bond P is a premium bond with a coupon rate of 8 percent, which makes annual payments, has YTM of 5 percent, and has ten years to maturity. If interest rates remain unchanged, what is the expected capital gains yield over the next year for Bond P?

1.45%

6.5%

3.55%

-1.5%
Answer #1

Let us assume the face value to be $100

Coupon = 8% *100 = 8

YTM = 5%

N = 10

Calculate the price of bond today (PV)

PV = Coupon * PVIFA (n,i) + FV * PVIF (n,i)

= 8 * PVIFA (10 ,5%) + 100 * PVIF (10, 5%)

= 8 *7.7217 + 100 * 0.6139

= 123.16

After one year we need to calculate the bond price with all parameters same except N = 9

PV = Coupon * PVIFA (n,i) + FV * PVIF (n,i)

= 8 * PVIFA (9 ,5%) + 100 * PVIF (9, 5%)

= 8 *7.1078 + 100 * 0.6446

= 121.32

One year capital gain yield = (121.32 - 123.16) /123.16

= - 1.50%